# How does the optimizer implicitly bias the model merging loss landscape?

### v1

@@ 3.4 Batch size and data augmentation
Figure 4: Batch size and data augmentation control the noise during the optimization dynamics. (left) A smaller batch size has more noise and improves the merging effectiveness. (right) Data augmentation improves performance and retains merging properties.
@@
We now consider two additional components that affect the effective noise scale: batch size and data augmentation. For batch size, we use the same setup as in Section 3.2, except that we sweep across different batch sizes and fix the training steps to 200k steps instead of using epochs. For data augmentation, we simply turn off the augmentation during the training phase. The experimental setup uses CIFAR100 with Resnet18. Further results can be found in Section B.1 and Section B.2.
@@
Batch size. The stochastic gradient of a minibatch $\hat{g}$ is unbiased but its variance scales as $Var(\hat{g})\propto\sigma^{2}/B$, where $\sigma^{2}$ is the per-sample variance and $B$ the batch size. Prior work has emphasized that this inverse scaling underlies the implicit regularization of SGD: smaller batch sizes inject more gradient noise, often leading to flatter solutions and better generalization (Jastrzebski et al., 2017; Smith & Le, 2018; Keskar et al., 2016). Our results in Figure 4 (left) extend this observation to model merging. We find that solutions obtained with smaller batch sizes are more compatible under linear interpolation: the smallest setup with $B=16$ achieves a median accuracy gain of $1?\%$, while a larger setup with $B=256$ yields almost no benefit. Thus, in addition to generalization, batch-size–induced noise also improves the merging effectiveness of independently trained models.
@@
Data augmentation. Data augmentation can likewise be viewed as injecting stochasticity into the optimization process: by applying random transformations to the data, the effective gradient covariance $\Sigma_{\mathcal{A}}$ changes, introducing additional variance beyond minibatch sampling. Previous work has argued that augmentation acts as a form of implicit regularization and invariance enforcement (Hernández-García & König, 2018; Yun et al., 2019), with recent perspectives interpreting augmentation as an additional source of optimization noise (Hanin & Sun, 2021). Our results in Figure 4 (right) show that augmentation not only improves single-model accuracy but also retains merging effectiveness. Interestingly, even without augmentation, a sufficiently large learning rate can yield positive merging gains, though this effect is not universal (Figure 10). Overall, augmentation-induced noise complements the minibatch noise and learning-rate noise, shaping solutions that are both stronger individually and compatible when merged.
@@ 3.5 What about language modeling?
Figure 5: Larger learning rate and weight decay enable more effective merging in language modeling. (left) A larger learning rate has a better loss gain. (right) Adding a larger weight decay offers further merging gains. Section B.4 shows the scatter plots.
Now we consider a language modeling task using the TinyStories dataset (Eldan & Li, 2023). We train a small GPT Transformer model with two layers using the AdamW optimizer with a constant learning rate for 200k steps and save a checkpoint every 2k steps, following the setup at Section A.1. Two endpoint models are trained for an additional 20k steps using a decayed learning rate scheduler. We use the loss performance gain to quantify the merging process. The lower the loss gain, the easier the merging process.
The results in Figure 5 extend our previous findings to the language domain. For the learning rate experimental setup, we fix $\text{wd}=0$. Figure 13 (left) shows that a larger learning rate, such as $\text{lr}=0.001$, requires fewer steps to reach a loss value of $2.20$ compared to $\text{lr}=0.0001$. Not only that, Figure 5 shows that a larger learning rate also has the implicit bias of simplifying the merging process, as measured by a lower loss gain. This behaviour is similar to the results from vision in Section 3.2. When weight decay regularization is added to the equation, there are further merging benefits. We fix the $\text{lr}=0.001$ and sweep across weight decay. Figure 5 (right) shows that a larger weight decay leads to a better loss gain than smaller ones. Specifically, the largest weight decay, such as $\text{wd}=0.1$, has the best loss gain, but also has a slower convergence ($x$-axis). The second largest weight decay of $\text{wd}=0.01$ has a similar convergence speed as the smaller one, in addition to better loss gain. Lastly, smaller values have similar results to training without weight decay.
@@ 3.6 What about transfer learning?
Figure 6: Merging effectiveness in transfer learning for ID and OOD data. (top) Accuracy gain linearly correlates with learning rate measured via Pearson correlation $r=0.981$. (bottom) However, a larger learning rate leads to a suboptimal merged model, despite having the largest accuracy gain.
In the previous sections, we analyzed settings where models are trained from scratch. Now we consider setups involving transfer learning. We finetuned only the vision encoder of the pretrained model CLIP ViT-B/16 (Radford et al., 2021) on the WILDS-FMoW (Koh et al., 2021) dataset using the AdamW optimizer with cosine scheduler. Since varying the learning rate changes the speed of convergence, we carefully tune the number of training epochs for each setup and ensure proper convergence (details in Section A.2). We train three seeds for each setup and merge each different pair, obtaining three different merged models per learning rate.
The results on the left of the Figure 6 shows that a larger learning rate identifies solutions that are easier to merge. Specifically, the smallest values lie in a flatter loss landscape region where the performance gain is $4\times$ smaller than the largest learning rate when merged. The Pearson correlation coefficient is $r=0.981$, indicating an almost perfect linear correlation between accuracy gain and learning rate. Note, however, that one should not blindly use the largest learning rate. Figure 6 on the right shows that the merged models with the best performance are the one with a moderate learning rate, as also observed by (Wortsman et al., 2022). The largest learning rate setup has the largest accuracy gain, but the worst-performing single model. Section B.3 presents similar results using different datasets and a pretrained model.
@@ 4 The optimizer’s implicit bias on task arithmetic
In the previous section, we have seen how the optimizer implicitly biases the loss landscape of linear interpolation merging. We now consider task arithmetic interpolation, which defines a different subspace of solutions. This section studies how the principal optimizer component, the learning rate, affects the loss landscape of task arithmetic merging.
@@ 4.1 Loss landscape of task arithmetic
(a) w/ Pretrained weight
(b) w/o Pretrained weight
Figure 7: Task arithmetic loss landscape drastically changes depending on the initialization model. (a) With a pretrained initialization on ImageNet, larger learning rate solutions have higher gain and are more robust to task arithmetic interpolation. (b) Without a pretrained weight, a larger learning rate solution lies in a sharper minima (i.e. more sensitive to $\alpha$ changes).
So far, we have only considered merging using linear interpolation (see Equation 1). Task arithmetic interpolates two models along a different subspace compared to linear interpolation, identifying functionally different solutions. We apply task arithmetic interpolation to two settings:
  1. (a)
    Models w/ pretraining weight from Section 3.6 (i.e. pretraining dataset is different from the finetuning dataset). Task arithmetic is applied to a base model and two task vectors. The base model $\bm{\theta}_{base}$ is the pretrained model CLIP ViT-B/16, and the task vectors are the finetuned models with different random seeds.
  2. (b)
    Models w/o pretraining weight from Section 3.2 (i.e. pretraining dataset is the same as the finetuning dataset). Task arithmetic is applied to a base model and two task vectors. For each checkpoint, the base model $\bm{\theta}_{base}$ is the checkpoint itself, and the task vectors are obtained from the endpoint models $\tau_{A}=\bm{\theta}_{A}-\bm{\theta}_{base}$ and $\tau_{B}=\bm{\theta}_{B}-\bm{\theta}_{base}$.
For each learning rate setup, we traverse the subspace defined by the task arithmetic interpolation by changing the coefficient $\alpha$. This measures the performance change as a function of $\alpha$, which can also be seen as a measure of landscape flatness. For simplicity, we use the same $\alpha$ for the two task vectors when applying task arithmetic.
The results in Figure 7 show the robustness of each learning rate to task arithmetic interpolation for CIFAR10. There is a clear dichotomy between the two settings. In setting (a), a larger learning rate identifies merged solutions that are more robust to $\alpha$-interpolation, corresponding to a flatter landscape (Andriushchenko et al., 2023a). However, in setting (b), the opposite is true. This highlights that a larger learning rate has to be used together with a suitable initialization to achieve a smoother and flatter landscape (Wortsman et al., 2022). Furthermore, as in linear interpolation merging, a “too large” learning rate becomes unstable. Section C.3 presents further results for different datasets.
@@ 4.2 Loss landscape of merging different tasks
Figure 8: Task arithmetic merging of models trained on two different tasks. (left) The merged models are finetuned using the same learning rate. Larger learning rate solutions have better performance and are more robust to task arithmetic interpolation, unless it is too large. (right) The models are merged using different learning rates. Merging pairs of similar and larger learning rates has the best performance. Results are averaged over three seeds.
@@
We now consider task arithmetic merging of two models sharing the same initialization $\bm{\theta}_{base}$ finetuned on two different, but similar, tasks $t$. As in Section 3.6, we finetune one CLIP ViT-B/16 on task $t_{1}$ WILDS-FMoW and another on task $t_{2}$ RESISC45 (Cheng et al., 2017). Then, task arithmetic merging is applied to merge the two models. To quantify the merging success, we use the averaged normalized accuracy, which measures the average ratio of the merged model performance over each single model performance (details in Section A.3).
@@
The results in Figure 8 show how the learning rate affects the merging effectiveness of two models trained on two different tasks. In the Figure 8 on the left, as a proxy of the task arithmetic loss landscape, we merge and study the robustness of models finetuned using the same hyperparameters when interpolating $\alpha$. We observe that the larger learning rate solutions perform better compared to the smaller ones (except for $\text{lr}=0.0001$, which is the limit for stability). Moreover, larger values are also more robust to changes of $\alpha$, representing flatter minima connecting the two different tasks. On the right of Figure 8, we merge models finetuned with different hyperparameters. The merged models with the best performance are those merged with similar and moderately large learning rates (near the antidiagonal). Merging models with a larger learning rate can result in better performance, but at the cost of losing flexibility for merging with other configurations. In particular, the largest learning rate $\text{lr}=0.0001$ is the most unstable to merge with different learning rate models. Ilharco et al. (2023) also observed performance degradation when merging models trained with too large learning rates. Section C.4 reports additional results with further $\alpha$ values.
5 Related works
@@
Model merging. Early works on merging independently trained solutions on the same task can be found on mode connectivity (Garipov et al., 2018; Draxler et al., 2018). Linear mode connectivity has a stricter condition such that connecting paths are linear (Frankle et al., 2020; Neyshabur et al., 2020). When this is not possible, re-basin methods can be used to reparametrize the solution and restore the linear connectivity (Entezari et al., 2022; Ainsworth et al., 2023; Theus et al., 2025). Built upon these results, model merging methods have been developed to increase the performance on a single task (Wortsman et al., 2022) or to combine models trained on different tasks into one (Matena & Raffel, 2022; Ilharco et al., 2023)Yadav et al. (2025) provides a comprehensive survey of the latest merging methods.
@@
Optimization dynamics. Standard optimization theory (Garrigos & Gower, 2023) shows that both batch sizes and learning rates drastically affect stability and convergence properties of SGD. In particular, through an analysis of SGD’s stationary distribution on simple quadratic potentials (Jastrzebski et al., 2017), it is possible to evince that, for single model training, the loss statistics at convergence only depend on the ratio between batch size and learning rates – as also validated empirically by Smith et al. (2020). In turn, either high learning rates or low batch sizes are known to favor flat minima (Keskar et al., 2016). While for more sophisticated optimizers, correlations between batch size, learning rates, and generalization might be more complex (Zhang et al., 2019; Malladi et al., 2022), other factors might more severely affect simple relations, such as non-Gaussianity (Simsekli et al., 2019) of gradient noise and non-convexity (Xie et al., 2021).

### v2

@@ 3.4 Batch size, momentum, and data augmentation
Figure 4: Batch size, momentum, and data augmentation control the noise during the optimization dynamics. (a, b) A small $B$ or a larger $\mu$ has more noise and improves the merging effectiveness. (c) Data augmentation improves performance and merging effectiveness.
@@
We consider three additional components affecting the effective noise scale: batch size $B$, momentum $\mu$, and data augmentation $\Sigma_{\mathcal{A}}$. We train ResNet18 on CIFAR100 with SGD. For batch size, we use the same setup as in Section 3.2, except that we sweep across different $B$ and fix the training steps to 200k steps instead of epochs. Similarly, we sweep across different momentum values. For augmentation, we simply turn off the augmentation during the training. Additional results on different datasets and scheduler choices are in Sections E.1, B.1 and B.2.
@@
Batch size. Stochastic gradient of a minibatch $\hat{g}$ is unbiased but its variance scales as $Var(\hat{g})\propto\sigma^{2}/B$, where $\sigma^{2}$ is the per-sample variance and $B$ the batch size. Prior work has emphasized that this inverse scaling underlies the implicit bias of SGD: smaller batch sizes inject more gradient noise, leading to flatter solutions and better generalization (Jastrzebski et al., 2017; Smith and Le, 2018; Keskar et al., 2016)Figure 4 (a) extends this observation to model merging. Solutions obtained with a smaller $B$ are more compatible under merging. The smallest $B=16$ achieves a median accuracy gain of $1\%$, while a larger setup with $B=256$ has almost no benefit. This suggests that $B$ induced noise can improve the merging effectiveness.
Momentum. Optimizers use momentum $\mu$ to accumulate an exponentially weighted moving average of past gradients. Traditionally understood as an acceleration mechanism for escaping shallow local minima and traversing flat regions (Polyak, 1964; Sutskever et al., 2013), momentum also alters the effective noise characteristics of SGD. Figure 4 (b) shows that models trained with a larger $\mu=0.9$ exhibit consistently better mergeability than those trained with a lower or no $\mu$, achieving median accuracy gains of $1.0\%$ compared to $0.2\%$ gains for low damped trajectories. Together with batch size, larger momentum shapes a better landscape for model merging.
@@
Augmentation. Data augmentation can also be viewed as injecting stochasticity into the optimization process. By applying random transformations to the input, the gradient covariance $\Sigma_{\mathcal{A}}$ incorporates additional variance beyond minibatch sampling. Previous work interprets augmentation as a form of implicit regularization (Hernández-García and König, 2018), or as an additional source of optimization noise (Hanin and Sun, 2021). Figure 4 (c) show that augmentation not only improves single-model accuracy but also retains merging effectiveness. And even without augmentation, a larger learning rate can yield positive merging gains. Augmentation induced noise complements the minibatch and learning rate noises, shaping the local and global loss landscape of merged models.
@@ 3.5 What about language modeling?
Figure 5: Larger learning rate and weight decay enable more effective merging in language modeling. (top) A larger learning rate $\eta$ achieves a better loss gain. (bottom) Adding a larger $\lambda$ offers further merging gains.
Now we consider a language modeling task using the TinyStories dataset (Eldan and Li, 2023). We train a small GPT Transformer model with two layers using the AdamW optimizer with a constant learning rate for 200k steps and save a checkpoint every 2k steps, following the setup at Section A.1. Two endpoint models are trained for an additional 20k steps using a decayed learning rate scheduler. We use the loss performance gain to quantify the merging effectiveness.
Figure 5 extends and supports our results from the previous sections to the language domain. Starting from the learning rate $\eta$ experiments, we fix $\lambda=0$. Figure 5 shows that a larger learning rate simplifies the merging process as measured by a lower loss gain, suggesting an implicit bias effect on the loss landscape. Additionally, Appendix Figure 13 (top) shows that a larger learning rate $\eta=1e-3$ converges faster to a loss value of $2.20$ compared to $\eta=1e-4$. These observations match the previous results from the vision domain. When adding weight decay $\lambda$ to the equation, further merging benefits is observed. We fix the $\eta=1e-3$ and sweep across $\lambda$ values. Figure 5 (bottom) shows that a larger $\lambda$ leads to a better loss gain than a smaller value. The largest weight decay, such as $\lambda=1e-1$, has the best loss gain, but also has a slower convergence (Figure 13). The second largest weight decay of $\lambda=1e-2$ has a similar convergence speed as $\lambda=1e-3$, but with better loss gain. Lastly, the smallest values have negligible effect as they have similar results as training without weight decay.
@@ 3.6 What about transfer learning?
Figure 6: Merging effectiveness in transfer learning for ID and OOD data. (top) Accuracy gain linearly correlates with learning rate. (bottom) However, a larger learning rate leads to a suboptimal merged model, despite having the largest accuracy gain.
The setups of previous sections train and bifurcate models on a single task, performing model merging between the bifurcated endpoints at the end. Now we consider transfer learning, where the pretraining and finetuning tasks differ. We use the vision encoder CLIP ViT-B/16 (Radford et al., 2021) pretrained on ImageNet1K and finetune it on the WILDS-FMoW (Koh et al., 2021) dataset using AdamW with cosine scheduler. Since varying the learning rate changes the speed of convergence, we tune the number of training epochs for each setup and ensure convergence to near-zero training loss (details in Section A.2). We train three seeds per configuration and merge each pair, resulting in three merged models per $\eta$. Additional results using the SGD optimizer instead of AdamW are in Section E.2.
Figure 6 (top) shows that a larger learning rate $\eta$ identifies solutions that are easier to merge. The smallest $\eta=1e-6$ lie in a flatter loss landscape region where the performance gain is $4\times$ smaller than the largest $\eta=1e-4$ when merged. The Pearson correlation coefficient is $r=0.981$, indicating an almost perfect linear correlation between accuracy gain and learning rate $\eta$. Note, however, that one should not blindly use the largest learning rate for finetuning. Figure 6 (bottom) shows that the merged models with the best performance are the ones with a moderate learning rate $\eta=3e-5$, as also observed by (Wortsman et al., 2022). The largest learning rate setup has the largest accuracy gain, but the worst-performing single model. Section B.3 presents similar results using different datasets and a pretrained model.
@@ 4 The optimizer’s implicit bias on task arithmetic
In the previous section, we have seen how the optimizer implicitly biases the loss landscape for linear interpolation merging. We now turn to task arithmetic (TA) interpolation, which defines a different subspace of solutions compared to the previous method.
@@ 4.1 Loss landscape of task arithmetic
(a) Transfer learning
(b) Non-transfer learning
Figure 7: Task arithmetic loss landscape based on the model initialization. (a) In transfer learning with a pretrained initialization on ImageNet1K, larger learning rate $\eta$ solutions have a larger accuracy gain and a smoother loss landscape (i.e. more robust to $\alpha$-interpolation). (b) In non-transfer learning however, a larger $\eta$ solutions lie in a sharper minima.
Task arithmetic interpolates two models along a different subspace compared to linear interpolation merging, identifying functionally different solutions. Here, we apply task arithmetic merging to study how it implicitly bias the loss landscape in two different settings: (a) Transfer learning setup. The pretraining and finetuning datasets do not overlap, same as in Section 3.6. Task arithmetic is applied using the pretrained base model $\bm{\theta}_{base}$ (ConvNext-Tiny pretrained on ImageNet1K) and the finetuned models on CIFAR10; (b) Non-transfer learning setup. The pretraining task is the same as the finetuning dataset, same as in Section 3.2. Task arithmetic is applied using the model before the bifurcation as the base model $\bm{\theta}_{base}$ and the endpoint models, all trained on CIFAR10. For each configuration, we change the $\alpha$-interpolation coefficient to traverse the subspace defined by the task arithmetic merging. This measures the performance change as a function of $\alpha$, which captures the shape of the loss landscape geometry. For simplicity, we fix $\alpha=0.5$ for the two task vectors when using task arithmetic merging.
Figure 7 shows the robustness of each configuration to task arithmetic interpolation, highlighting a dichotomy between the two settings. In (a), a larger learning rate identifies merged solutions that are more robust to $\alpha$-interpolation changes, corresponding to a flatter loss landscape (Andriushchenko et al., 2023a). While in (b), the opposite is observed, that is a smaller $\eta$ is associated to a flatter surface. These results highlight the importance of the $\bm{\theta}_{base}$. A larger learning rate should pair with a suitable initialization to shape a smoother and flatter loss landscape (Wortsman et al., 2022). Lastly, as in linear interpolation merging, we observe that a “too large” learning rate becomes unstable in both settings (a) and (b). Section C.3 presents additional results for different datasets.
@@ 4.2 Loss landscape of merging different tasks
(a) Loss landscape $\alpha$-interpolation
(b) Normalized accuracy cross-$\eta$
Figure 8: Merging two models trained on different tasks using TA or TIES merging. (a) Models trained using the same $\eta$ are merged. Larger $\eta$ solutions are more robust to $\alpha$-interpolation, up to the critical point $\eta=3e-5$. TIES shapes a wider loss landscape than TA. (b) Merging models trained with different $\eta$. Merging pairs of similar, larger $\eta$ yields the best performance.
@@
We now consider merging two models sharing the same initialization $\bm{\theta}_{base}$ finetuned on two different tasks $t$. We finetune one $\bm{\theta}_{base}$ model CLIP ViT-B/16 on $t_{1}$ WILDS-FMoW and another on $t_{2}$ RESISC45 (Cheng et al., 2017). Then, TA/TIES (Yadav et al., 2023) is applied to merge the two models. We measure the loss landscape geometry via $\alpha$-interpolation. And to quantify the merging success, we use the averaged normalized accuracy, which averages the ratio of the merged model performance over each single model performance (defined in Section A.3). Further details on the hyperparameters are in Section A.2. Results are averaged over three random seeds.
@@
Figure 8 shows how the primary optimizer component ($\eta$) impacts the merging effectiveness of models trained on different tasks. Figure 8 (a) presents the loss landscape geometry of the subspace defined by TA/TIES. We study the robustness of models trained using the same hyperparameters when interpolating $\alpha$. We observe that training with a larger learning $\eta$ yields better performance compared to smaller ones in both merging methods (except for $\eta=1e-4$, which is the upper limit for stability). Moreover, larger $\eta$ are also more robust to changes of $\alpha$, corresponding to a flatter minima solution. In the Figure 8 (b), merged models are trained with different learning rates $\eta$, and we present the best $\alpha$ per merging method. The highest performance merged models are those combined with similar and moderately larger $\eta$ (across the antidiagonal). Note that merging task vectors from increasingly larger $\eta$ have worse compatibility when merging with other $\eta$ models. For example, the largest learning rate $\eta=1e-4$ is the most incompatible to merge with all the others. Ilharco et al. (2023) also observed performance degradation when merging models trained with too large $\eta$. Our results show that merging two moderately larger $\eta$ models have the highest accuracy but lose the compatibility property.
Lastly, the results in Figure 8 (b) also suggests that TIES merging can better counteract the noise induced by larger $\eta$, yielding a $2\%$ improvement compared to task arithmetic at the best setup $\eta=3e-5$ ($88.0\%$ vs $85.9\%$). In the loss landscape analysis, TIES also shapes a wider loss landscape compared to TA merging as it is more stable across $\alpha$-interpolation. Appendix F contains similar results with TIES merging models trained on three different tasks.
5 Permutation symmetry and feature alignment
We show how the effective noise can shape the geometry of a minimum under permutation symmetry, and that merging requires complementary features to achieve performance gains.
Permutation symmetry. Prior work conjectures that two models $\theta_{A}$ and $\theta_{B}$ trained on the same task from different random initializations are equivalent up to permutation symmetries (Entezari et al., 2022; Ainsworth et al., 2023). Accordingly, there may exist a permutation $\pi$ such that the rebased model $\pi(\theta_{A})$ aligns with $\theta_{B}$, enabling linear mode connectivity. In our experiments, larger effective noise simplifies this rebasing process. Following Section 3.2, we train two ResNet18 models on CIFAR100 with independent random initializations and rebase one model via weight-based matching (Ainsworth et al., 2023). To visualize the effect of rebasing, we construct a 2D slice of the loss landscape by fixing three points as in Izmailov et al. (2018). Figure 9 (left, middle) shows that larger $\eta$ corresponds to a wider minimum and a flatter interpolation path between $\theta_{B}$ and $\pi(\theta_{A})$, indicating that the solutions are more easily alignable under permutation. We observe a similar phenomenon for Transformer language models (see Appendix D.1).
Feature alignment. To connect our effective noise results to representation level behavior, we measure feature alignment between the two branched checkpoints using linear CKA (Kornblith et al., 2019) on penultimate-layer activations (2048 test samples), and relate it to the accuracy gain from merging. As shown in Figure 9 (right), increasing the effective noise (here induced by larger $\eta$) simultaneously increases merge gains and decreases feature alignment, indicating more diverse features across branches. Together with the non-monotonic dependence of mergeability on effective noise, this suggests that the intermediate/critical noise regime promotes the kind of feature diversity that merging can exploit, whereas low-noise training produces highly aligned representations with little merging benefits.
Figure 9: Larger learning rate $\eta$ identifies a (left, middle) broader basin that simplifies model permutation, and (right) more diverse features that achieve higher accuracy gain. The setup uses ResNet18 trained on CIFAR100 as in Section 3.
6 Related work
@@
Model merging. Early works on merging independently trained solutions on the same task can be found on mode connectivity (Garipov et al., 2018; Draxler et al., 2018). Linear mode connectivity has a stricter condition such that connecting paths are linear (Frankle et al., 2020; Neyshabur et al., 2020). When this is not possible, re-basin methods can be used to reparametrize the solution and restore the linear connectivity (Entezari et al., 2022; Ainsworth et al., 2023; Theus et al., 2025). Built upon these results, model merging methods have been developed to increase the performance on a single task (Wortsman et al., 2022) or to combine models trained on different tasks into one (Matena and Raffel, 2022; Ilharco et al., 2023). We refer the reader to Yadav et al. (2025), which provides a comprehensive survey of the latest merging methods.
@@
Optimization dynamics. Standard optimization theory (Garrigos and Gower, 2023) shows that both batch sizes and learning rates drastically affect stability and convergence properties of SGD. In particular, through an analysis of SGD’s stationary distribution on simple quadratic potentials (Jastrzebski et al., 2017), it is possible to evince that, for single model training, the loss statistics at convergence only depend on the ratio between batch size and learning rates – as also validated empirically by Smith et al. (2020). In turn, either high learning rates or low batch sizes are known to favor flat minima (Keskar et al., 2016). While for more sophisticated optimizers, correlations between batch size, learning rates, and generalization might be more complex (Zhang et al., 2019; Malladi et al., 2022), other factors might more severely affect simple relations, such as non-Gaussianity (Simsekli et al., 2019) of gradient noise and non-convexity (Xie et al., 2021).